# Counting points on generic character varieties

### v2

@@ 7.7. Types for $G=G_{2}$
$[G_{2},\phi_{1,0}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,1}]$ $6q^{5}\Phi_{1}^{2}\Phi_{6}$ $2^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q\Phi_{2}^{2}\Phi_{3}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,2}]$ $2q^{5}\Phi_{1}^{2}\Phi_{3}$ $2^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q\Phi_{2}^{2}\Phi_{6}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,6}]$ $\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}$ $1$ $12$ $1$ $1$ $1$
$[{A_{2}},3^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $1$ $1$ $6$ $1$ $3$ $2$
$[{A_{2}},2^{1}1^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{3}$ $4^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q\Phi_{2}$ $2$ $6$ $1$ $3$ $2$
$[{A_{2}},1^{3}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q^{3}$ $1$ $6$ $1$ $3$ $2$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 2^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 1^{2}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-2$
$[T,1^{1}]$ $q^{6}\Phi_{1}^{2}$ $12^{n}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $12$
@@ 7.8. Types for $\mathfrak{g}=\mathfrak{g}_{2}$
$[G_{2},0]$ $q^{14}$ $q^{6n}\Phi_{2}^{2n}\Phi_{3}^{n}\Phi_{6}^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $12$ $1$ $1$
$[G_{2},A_{1}]$ $q^{8}$ $q^{6n}\Phi_{1}\Phi_{2}^{n+1}\Phi_{3}^{n+1}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}\Phi_{2}\Phi_{3}\Phi_{6}$ $\Phi_{2}\Phi_{3}$ $12$ $1$ $1$
$[G_{2},\tilde{A_{1}}]$ $q^{6}$ $q^{6n+2}\Phi_{1}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}(2q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{2}\Phi_{1}\Phi_{2}\Phi_{3}\Phi_{6}$ $(2q+1)\Phi_{2}$ $12$ $1$ $1$
$[G_{2},G_{2}(a_{1})_{1^{3}}]$ $q^{4}$ $\frac{1}{6}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}(4q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $14q+$ 1 $12$ $1$ $1$
$[G_{2},G_{2}(a_{1})_{2^{1}1^{1}}]$ $q^{4}$ $\frac{1}{2}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}(2q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $12q+$ 1 $12$ $1$ $1$
$[G_{2},G_{2}(a_{1})_{3^{1}}]$ $q^{4}$ $\frac{1}{3}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{n+2}\Phi_{3}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{2}$ $12$ $1$ $1$
$[G_{2},G_{2}]$ $q^{2}$ $q^{6n+4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$
$[A_{1},1^{2}]$ $q^{4}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $3$ $-1$
$[A_{1},2^{1}]$ $q^{2}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{1}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $3$ $-1$
$[A_{1}^{\prime},1^{2}]$ $q^{4}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $3$ $-1$
$[A_{1}^{\prime},2^{1}]$ $q^{2}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{1}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $3$ $-1$
$[T,0]$ $q^{2}$ $5\cdot 12^{n-1}q^{6n+6}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $5$
@@ 7.10. Counting polynomials when $(g,n)=(0,3)$
$B_{3}$ $22q^{12}+18q^{11}+88q^{10}+312q^{9}+952q^{8}+2478q^{7}+4980q^{6}+\cdots+18q+$ 2
$C_{3}$ $22q^{12}+18q^{11}+88q^{10}+320q^{9}+970q^{8}+2506q^{7}+5060q^{6}+\cdots+18q+$ 2
$D_{4}$ $4q^{16}+48q^{15}+308q^{14}+1408q^{13}+5140q^{12}+16176q^{11}$ $+43748q^{10}+96256q^{9}+152864q^{8}+\cdots+48q+4$
$F_{4}$ $q^{40}+12q^{39}+77q^{38}+352q^{37}+1287q^{36}+4004q^{35}+11010q^{34}+27444q^{33}$ $+63155q^{32}+136096q^{31}+277849q^{30}+542780q^{29}+1023705q^{28}+1879344q^{27}$ $+3384741q^{26}+6009812q^{25}+10498044q^{24}+17873468q^{23}+29224321q^{22}$ $+44790488q^{21}+58508548q^{20}+\cdots+12q+1$
$E_{6}$ $3q^{60}+54q^{59}+510q^{58}+3366q^{57}+17442q^{56}$ $+75579q^{55}+284829q^{54}+958746q^{53}$ $+2938005q^{52}+8314473q^{51}+21970542q^{50}$ $+54685197q^{49}+129125661q^{48}+290973411q^{47}$ $+628951404q^{46}+1310014803q^{45}+2640078312q^{44}$ $+5167269447q^{43}+9854714367q^{42}+18363440685q^{41}$ $+33501157887q^{40}+59902887102q^{39}+104998780206q^{38}+180252368460q^{37}$ $+302421156681q^{36}+493955088993q^{35}+780185905473q^{34}+1178431874241q^{33}$ $+1671208058577q^{32}+2158923553011q^{31}+2408708043594q^{30}+\cdots+510q^{2}+54q+3$

### v3

@@ 7.7. Types for $G=G_{2}$
$G_{2},\phi_{1,0}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ 11 $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ 11 11 $12$ 11 11 11
$G_{2},\phi_{1,3}^{\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ 11 $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ 11 $12$ 11 11 11
$G_{2},\phi_{1,3}^{\prime\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ 11 $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ 11 $12$ 11 11 11
$G_{2},\phi_{2,1}]$ $6q^{5}\Phi_{1}^{2}\Phi_{6}$ $2^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q\Phi_{2}^{2}\Phi_{3}$ 22 $12$ 11 11 11
$G_{2},\phi_{2,2}]$ $2q^{5}\Phi_{1}^{2}\Phi_{3}$ $2^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q\Phi_{2}^{2}\Phi_{6}$ 22 $12$ 11 11 11
$G_{2},\phi_{1,6}]$ $\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ 11 $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}$ 11 $12$ 11 11 11
${A_{2}},3^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ 11 11 66 11 33 22
${A_{2}},2^{1}1^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{3}$ $4^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q\Phi_{2}$ 22 66 11 33 22
${A_{2}},1^{3}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q^{3}$ 11 66 11 33 22
${A_{1}\times A_{1}^{\prime}},2^{1}\otimes 2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ 11 11 44 33 22 11
${A_{1}\times A_{1}^{\prime}},2^{1}\otimes 1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $^{2}\Phi_{1}^{2}\Phi_{2}^{2} qq$ 11 44 33 22 11
${A_{1}\times A_{1}^{\prime}},1^{2}\otimes 2^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $^{2}\Phi_{1}^{2}\Phi_{2}^{2} qq$ 11 44 33 22 11
${A_{1}\times A_{1}^{\prime}},1^{2}\otimes 1^{2}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ 11 44 33 22 11
${A_{1}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $2\cdot 6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 11 22 33 11 4-4
${A_{1}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $2\cdot 6^{n}$ $\Phi_{1}^{2}\Phi_{2} qq$ 11 22 33 11 4-4
${A_{1}^{\prime}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 11 22 33 11 2-2
${A_{1}^{\prime}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $6^{n}$ $\Phi_{1}^{2}\Phi_{2} qq$ 11 22 33 11 2-2
$T,1^{1}]$ $q^{6}\Phi_{1}^{2}$ $12^{n}$ $\Phi_{1}^{2}$ 11 11 11 11 11 $12$
@@ 7.8. Types for $\mathfrak{g}=\mathfrak{g}_{2}$
$G_{2},0]$ $q^{14}$ $q^{6n}\Phi_{2}^{2n}\Phi_{3}^{n}\Phi_{6}^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ 11 $\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $12$ 11 11
$G_{2},A_{1}]$ $q^{8}$ $q^{6n}\Phi_{1}\Phi_{2}^{n+1}\Phi_{3}^{n+1}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}\Phi_{2}\Phi_{3}\Phi_{6}$ $\Phi_{2}\Phi_{3}$ $12$ 11 11
$G_{2},\tilde{A_{1}}$] $q^{6}$ $q^{6n+2}\Phi_{1}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}(2q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{2}\Phi_{1}\Phi_{2}\Phi_{3}\Phi_{6}$ $(2q+1)\Phi_{2}$ $12$ 11 11
$G_{2},G_{2}(a_{1})_{1^{3}}]$ $q^{4}$ $\frac{1}{6}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}(4q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $4q+1$ $12$ 11 11
$G_{2},G_{2}(a_{1})_{2^{1}1^{1}}]$ $q^{4}$ $\frac{1}{2}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}(2q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $2q+1$ $12$ 11 11
$G_{2},G_{2}(a_{1})_{3^{1}}]$ $q^{4}$ $\frac{1}{3}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{n+2}\Phi_{3}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{2}$ $12$ 11 11
$G_{2},G_{2}]$ $q^{2}$ $q^{6n+4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ 11 $12$ 11 11
$A_{1},1^{2}]$ $q^{4}$ $3\cdot 6^{n-1}q^{6n+5}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 $\Phi_{2}$ 22 33 1-1
$A_{1},2^{1}]$ $q^{2}$ $3\cdot 6^{n-1}q^{6n+5}\Phi_{1}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ 11 22 33 1-1
$A_{1}^{\prime},1^{2}]$ $q^{4}$ $3\cdot 6^{n-1}q^{6n+5}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ 11 $\Phi_{2}$ 22 33 1-1
$A_{1}^{\prime},2^{1}]$ $q^{2}$ $3\cdot 6^{n-1}q^{6n+5}\Phi_{1}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ 11 22 33 1-1
$[T,0]$ $q^{2}$ $5\cdot 12^{n-1}q^{6n+6}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ 11 11 11 11 55
@@ 7.10. Counting polynomials when $(g,n)=(0,3)$
$B_{3}$ $2q^{12}+18q^{11}+88q^{10}+312q^{9}+952q^{8}+2478q^{7}+4980q^{6}+\cdots+18q+2$
$C_{3}$ $2q^{12}+18q^{11}+88q^{10}+320q^{9}+970q^{8}+2506q^{7}+5060q^{6}+\cdots+18q+2$
$D_{4}$ $4q^{16}+48q^{15}+308q^{14}+1408q^{13}+5140q^{12}+16176q^{11}$ $43748q^{10}+96256q^{9}+152864q^{8}+\cdots+48q+4$
$F_{4}$ $q^{40}+12q^{39}+77q^{38}+352q^{37}+1287q^{36}+4004q^{35}+11010q^{34}+27444q^{33}$ $63155q^{32}+136096q^{31}+277849q^{30}+542780q^{29}+1023705q^{28}+1879344q^{27} +$$3384741q^{26}+6009812q^{25}+10498044q^{24}+17873468q^{23}+29224321q^{22} +$$44790488q^{21}+58508548q^{20}+\cdots+12q+1$
$E_{6}$ $3q^{60}+54q^{59}+510q^{58}+3366q^{57}+17442q^{56}$ $75579q^{55}+284829q^{54}+958746q^{53} +$$2938005q^{52}+8314473q^{51}+21970542q^{50} +$$54685197q^{49}+129125661q^{48}+290973411q^{47} +$$628951404q^{46}+1310014803q^{45}+2640078312q^{44} +$$5167269447q^{43}+9854714367q^{42}+18363440685q^{41} +$$33501157887q^{40}+59902887102q^{39}+104998780206q^{38}+180252368460q^{37} +$$302421156681q^{36}+493955088993q^{35}+780185905473q^{34}+1178431874241q^{33} +$$1671208058577q^{32}+2158923553011q^{31}+2408708043594q^{30}+\cdots+510q^{2}+54q+3$